# Basis Selection: Low-Rank Decomposition of Pretrained Large Language Models for Target Applications

### v1

@@
Abstract
@@ 1 Introduction
Our approach leverages the relationship between pretrained models and target applications. LLMs are typically pretrained on extensive datasets covering diverse tasks, which often share common elements. This shared pretraining generates synergies that contribute to the exceptional performance of LLMs. However, the differences between these tasks also introduce redundancy in the models. For a specific target application, we can remove these redundant components and retain only the useful parts of the LLMs. This approach allows us to compress the models while maintaining their performance on target applications as much as possible. Many scenarios only require supporting a certain type of application, making this approach effective in reducing the computing resource requirement and lowering inference costs.
@@
Figure 1: Key idea of Basis Selection: Identify the important bases for target applications and keep them during compression.
@@
However, how do we identify the beneficial and redundant components of LLMs for a specific application? In this work, we address this problem through the lens of matrix factorization. Singular Value Decomposition (SVD) [9] factorizes a weight matrix $W$ into the product of three matrices $U$, $S$, and $V$, i.e., $W=USV^{\mathrm{T}}=\sum_{i}{s_{i}\mathbf{u}_{i}{\mathbf{v}_{i}^{\mathrm{T}}}}$, where $s_{i}$ are (positive) singular values, and $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ are column vectors of $U$ and $V$ with unit norms. A recent study by Ferrando and Voita [8] demonstrates that these column vectors $u_{i}$ and $v_{i}$ carry specific meanings. For example, for an attention head typically activated in code generation tasks in the LLaMA 2-7B model [17], factorizing its weight matrix $W_{O}$ and de-embedding the resultant column vectors $\textbf{u}_{i}$ reveals tokens such as userId, groupId, filename, and className [8]. Obviously, these column vectors are highly useful for code generation tasks, but they are less beneficial for other tasks, such as mathematical reasoning.
@@
Inspired by this observation, we propose Basis Selection, a low-rank decomposition approach to effectively compress LLMs for target applications. Figure 1 illustrates the key idea of Basis Selection. We view each weight matrix in LLMs as a linear combination of bases $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$ with singular values $s_{i}$ as their weights. These bases are valuable representations stored in the pretrained model, learned from large pretraining datasets. For a target application, some bases are advantageous while many others are not. To select the bases beneficial for the target application, we propose retraining the singular values (i.e., the weights of the bases) while keeping the bases fixed, using the training set of the target application. After retraining, we prune the bases associated with small singular values, as they are less important for the target application, and retain only those with large singular values, which are most critical for the target application. This approach allows us to eliminate the redundant parts of the original LLMs and retain only the components essential for the target application. To handle the data distribution differences between the pretraining dataset and the target application, we also augment the model with new bases learned from the training set of the target application during the pruning process. This enables us to learn the new bases necessary for the target application that are absent in the pretrained model.
@@
We evaluate Basis Selection on two models—Llama 2-7B and Llama 2-13B [17]—and two tasks—mathematical reasoning and code generation. We evaluate the pass@1 accuracy of the compressed models on GSM8K [6] and MATH [10] for the mathematical reasoning task and HumanEval [3] and MBPP [2] for the code generation task. Compared to state-of-the-art baselines, our approach demonstrates significantly better performance when the compression ratio  [^1] The compression ratio is defined as the ratio of the original model size to the compressed model size. exceeds 6 for mathematical reasoning and 4 for code generation. This also implies that, for deep compression, our approach reduces the model size more significantly while maintaining similar accuracy to baseline methods.
@@
  • We propose Basis Selection, a low-rank decomposition approach to compress pretrained large language models for target applications. Basis Selection identifies the beneficial and redundant components of large language models by relearning the importance (i.e., singular values) of bases using the training set of the target application, and then selects bases based on their importance.
  • We evaluate Basis Selection across multiple tasks and models, demonstrating its superior performance in deep compression.
@@ 2 Related Work
Singular Value Decomposition (SVD) [9] has been applied to reduce the size of machine learning models. Prior research [18, 13, 7, 20, 15, 4, 1, 14] has developed various SVD algorithms to compress different components of models, such as DNNs, CNNs, and embedding layers for a range of applications including natural language processing, speech, and vision. The primary distinction between our work and these prior studies is that they do not relearn the importance of bases using the training data of target applications. Instead, they typically prune bases according to the singular values from the original or finetuned models. FWSVD [11] evaluates the importance of individual weights rather than bases during SVD. As highlighted in Sections 1 and 3, the bases hold significant physical meanings. Considering importance at this level of granularity results in improved performance. As shown in Section 4, our approach surpasses FWSVD in deep compression performance.
A recent study [16] applied SVD to large language models. Its focus is on determining the optimal rank for each layer, while our emphasis is on basis selection. The two methods are orthogonal but complementary and can be combined. Chen et al. [5] and Yu and Wu [19] suggest reconstructing bases based on feature mimicking. These approaches are orthogonal to ours—they concentrate on basis reconstruction, whereas we focus on basis selection. Their methods complement ours and can be integrated together to achieve enhanced compression results.
3 Basis Selection
In this section, we describe our proposed compression method, Basis Selection.
For a linear layer $\mathbf{y}=W\mathbf{x}+\mathbf{b}$, Singular Value Decomposition (SVD) factorizes its weight matrix $W\in\mathbb{R}^{n\times m}$ as the product of three matrices $U$, $S$, and $V$:
$W=USV^{\mathrm{T}}$ (1)
where $U=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $S=\mathrm{diag}\left(s_{1},\cdots,s_{r}\right)$, and $V=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$.
The values $\left\{s_{i}\in\mathbb{R},i=1,\cdots,r\right\}$ are positive singular values. [^2] We drop zero singular values and the corresponding columns of matrices $U$ and $V$. The vectors $\left\{\mathbf{u}_{i}\in\mathbb{R}^{n},i=1,\cdots,r\right\}$ and $\left\{\mathbf{v}_{i}\in\mathbb{R}^{m},i=1,\cdots,r\right\}$ are orthonormal, i.e., $\left\|\mathbf{u}_{i}\right\|=1$, $\left\|\mathbf{v}_{i}\right\|=1$, $\mathbf{u}_{i}\perp\mathbf{u}_{j}$, and $\mathbf{v}_{i}\perp\mathbf{v}_{j}$ if $i\neq j$.
@@ 3 Basis Selection
$W=\sum_{i=1}^{r}s_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$ (2)
Let matrix $W_{i}=\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$, then
$\left\|W_{i}\right\|=\sqrt{\mathrm{tr}\left(W_{i}^{\mathrm{T}}W_{i}\right)}=\sqrt{\mathrm{tr}\left(\mathbf{v}_{i}\mathbf{u}_{i}^{\mathrm{T}}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}\right)}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}\mathbf{v}_{i}^{\mathrm{T}})}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}^{\mathrm{T}}\mathbf{v}_{i})}=1$ (3)
$\left\langle W_{i},W_{j}\right\rangle=\mathrm{tr}\left(W_{i}^{\mathrm{T}}W_{j}\right)=\mathrm{tr}\left(\mathbf{v}_{i}\mathbf{u}_{i}^{\mathrm{T}}\mathbf{u}_{j}\mathbf{v}_{j}^{\mathrm{T}}\right)=0,\quad\text{if}\;i\neq j$ (4)
Therefore, $\left\{\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}},i=1,\cdots,r\right\}$ can be seen as a group of orthonormal bases in a subspace of $\mathbb{R}^{n\times m}$, and $\left\{s_{i},i=1,\cdots,r\right\}$ are their weights, making the weight matrix $W$ a linear combination of these bases.
@@
$\mathbf{y}=W\mathbf{x}+\mathbf{b}=\sum_{i=1}^{r}s_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}\mathbf{x}+\mathbf{b}=\sum_{i=1}^{r}s_{i}\left\langle\mathbf{x},\mathbf{v}_{i}\right\rangle\mathbf{u}_{i}+\mathbf{b}$ (5)
@@
In other words, for each basis (i.e., filter) $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$, the similarity between the input signal $\mathbf{x}$ and the unit direction vector $\mathbf{v}_{i}$ is measured by their inner product. This inner product is then multiplied by the (positive) singular value $s_{i}$ to determine the weight for the unit direction vector $\mathbf{u}_{i}$. The output signal $\mathbf{y}$ is the weighted sum of $\mathbf{u}_{i}$. Figure 2 illustrates this interpretation of the role of bases from the perspective of signal processing.
Figure 2: An interpretation of the role of bases from the perspective of signal processing.
Large language models pretrained on diverse datasets have bases that carry various meanings. Ferrando and Voita [8] investigate these meanings by factorizing the $W_{O}$ matrix of certain attention layers in Llama 2-7B and de-embedding the $\mathbf{u}$ vector associated with bases having large singular values. Table 3 presents key results from their study. It is observed that some bases have very specific meanings, such as those related to code generation. For target applications other than that, these bases are likely irrelevant. When using such bases as filters for irrelevant target applications, either the filter is not activated (due to a small inner product $\left\langle\mathbf{x},\mathbf{v}_{i}\right\rangle$), or worse, the filter is activated and harmful information is mixed into the output, leading to performance degradation. This suggests that pruning these bases can compress the models with negligible performance loss or, in some cases, even improve performance in the target application.
Table 1: The meaning of bases in Llama 2-7B (data are sourced from [8]).
Decoder Layer, Attention Head, Basis ID (sorted by singular values in descending order) Top ten most probable tokens corresponding to the vector $\mathbf{u}$ of the basis
Layer 17, Head 25, $W_{O}$ matrix, Basis 1 userID, groupId, filename, _className, artifactId, _filename, _Хронология, className, <s>, _javax
Layer 19, Head 17, $W_{O}$ matrix, Basis 4 _return, return, _Return, _Returns, _returning, Return, _returns, {CJK*}UTF8gbsn, _returned, returns
@@
In our approach, Basis Selection, we determine the importance of the bases from the pretrained model by retraining their singular values on the training set of the target application. The weight matrix $\widetilde{W}$ in Basis Selection is represented as:
$\widetilde{W}=\sum_{i=1}^{r}\tilde{{s}}_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}+\sum_{j=1}^{\tilde{r}}\tilde{\mathbf{u}}_{j}\tilde{\mathbf{v}}_{j}^{\mathrm{T}}$ (6)
In the first term, $\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$ represents the original bases in the pretrained model. To assess their importance, we initialize their weights $\tilde{s}_{i}$ with their original singular values and then retrain these weights (while keeping the bases fixed) on the training set of the target application. The aim is that, after retraining, the bases important for the target application will have larger singular values, whereas those that are useless or detrimental will have zero or very small singular values. This allows us to identify and prune the less useful bases. Relearning the importance of bases for the target application distinguishes our approach from previous methods. Prior approaches either use the singular values in the original model [18, 13, 7, 20, 15, 4, 1, 14, 16] or assess the importance of weight parameters, other than the importance of the bases, to prune them [11]. They do not relearn the importance of the bases specifically for the target application. From a signal processing perspective, this first term allows us to adjust the weight for each filter, catering to the needs of the target application.
In the second term, $\tilde{\mathbf{u}}_{j}$ and $\tilde{\mathbf{v}}_{j}$ are learnable vectors included for two primary purposes. First, due to differences in data distribution between the pretrained dataset and the target application, some bases necessary for the target application might be absent in the pretrained dataset. We use these vectors to learn such bases. Second, during pruning, although each pruned basis may individually have minimal impact on the target application, their cumulative performance loss can be significant. These additional vectors help compensate for the performance loss caused by the pruned bases. This second term is similar to LoRA [12]. The number of learnable vectors $\tilde{r}$ is referred to as the additional dimension. From a signal processing perspective, this second term allows us to include additional, new filters to enhance the model performance on the target application.
Algorithm 1 outlines our approach. Basis Selection takes a pretrained or fine-tuned model as input and gradually prunes bases in the original model. After each pruning step, it finetunes the learnable parameters $\tilde{s}_{i}$, $\tilde{\mathbf{u}}_{j}$, and $\tilde{\mathbf{v}}_{j}$ to offset the performance loss. Ultimately, a new weight matrix $\widetilde{W}$ with a smaller rank $r^{\prime}$ is learned. We perform a standard SVD on it, representing it as the product of matrices $U^{\prime}$, $S^{\prime}$, and $V^{\prime\mathrm{T}}$. We then replace the original layer with two new layers: $S^{\prime}V^{\prime\mathrm{T}}$ becomes the weight matrix of the first new layer, and $U^{\prime}$ becomes the weight matrix of the second new layer. This reduces the number of parameters from $nm$ to $(n+m)r^{\prime}$. The new model is subsequently further finetuned to enhance its performance on the target application.
@@
3 KeepRatioPerPruning = KeepRatio${}^{(1/\text{PruningTimes})}$;
4
5Convert the weight matrix of each layer in $M$ into the form of equation (6);
@@ 3 KeepRatioPerPruning = KeepRatio${}^{(1/\text{PruningTimes})}$;
7       Tune the learnable parameters in equation (6);
8

### v2

@@
11footnotetext: Equal contribution.22footnotetext: Corresponding author. Email: yangli1@iastate.edu
Basis Selection: Low-Rank Decomposition of Pretrained Large Language Models for Target Applications
Yang Li* yangli1@iastate.edu
Iowa State University / Meta
Daniel Agyei Asante* dasante@iastate.edu
Iowa State University
Changsheng Zhao cszhao@meta.com
Meta
Ernie Chang erniecyc@meta.com
Meta
Yangyang Shi yyshi@meta.com
Meta
Vikas Chandra vchandra@meta.com
Meta
Аннотация
@@ 1 Introduction
Our approach exploits the relationship between pretrained models and specific target applications. Large language models (LLMs) are typically pretrained on vast datasets encompassing a wide range of tasks, many of which share common characteristics. This shared pretraining fosters synergies that enhance the performance of LLMs. However, the diversity among these tasks also introduces redundancy into the models. As demonstrated by our interpretation results in Section 3 shows, LLMs contain a significant number of redundant components that are unnecessary for a specific target application. By removing these redundant parts and retaining only the relevant ones, we can reduce the model’s size while preserving its performance on the target application. Since many scenarios only require support for a specific type of application, this approach effectively lowers the computing resource requirements and reduces inference costs.
@@
Рис. 1: Basel: Identify and select the important bases for target applications during compression.
@@
However, how do we identify the beneficial and redundant components of LLMs for a specific application? In this work, we address this problem through the lens of matrix factorization. Singular Value Decomposition (SVD) (Golub & Van Loan, 1996) factorizes a weight matrix $\mathbf{W}$ into the product of three matrices $\mathbf{U}$, $\mathbf{S}$, and $\mathbf{V}$, i.e., $\mathbf{W}=\mathbf{USV}^{\top}=\sum_{i}{s_{i}\mathbf{u}_{i}{\mathbf{v}_{i}^{\top}}}$, where $s_{i}$ are (positive) singular values, and $\mathbf{u_{i}}$ and $\mathbf{v_{i}}$ are column vectors of $\mathbf{U}$ and $\mathbf{V}$ with unit norms. Our interpretation results show that these column vectors $\mathbf{u_{i}}$ and $\mathbf{v_{i}}$ may carry specific meanings. For instance, in the LLaMA 2-7B model (Touvron et al., 2023b), when factorizing the weight matrix $\mathbf{W_{O}^{h}W_{V}^{h}}$ of an attention head that is likely useful for code generation tasks, de-embedding the resulting column vectors $\mathbf{u_{i}}$ and $\mathbf{v_{i}}$ reveals tokens like _in, <0x0A>, _and, _to, and _for. These column vectors are evidently highly useful for code generation but may be less relevant for tasks such as mathematical reasoning.
@@
Inspired by this observation, we propose Basel, a low-rank decomposition approach to effectively compress LLMs for target applications. Figure 1 illustrates the key idea of Basel. We view each weight matrix in LLMs as a linear combination of bases $\mathbf{u_{i}}\mathbf{v_{i}}^{\mathrm{T}}$ with singular values $s_{i}$ as their weights. These bases are valuable representations stored in the pretrained model, learned from large pretraining datasets. For a target application, some bases are advantageous while many others are not. To select the bases beneficial for the target application, we propose retraining the singular values (i.e., the weights of the bases) while keeping the bases fixed, using the training set of the target application. After retraining, we prune the bases associated with small singular values, as they are less important for the target application, and retain only those with large singular values, which are most critical for the target application. This approach allows us to eliminate the redundant parts of the original LLMs and retain only the components essential for the target application. To handle the data distribution differences between the pretraining dataset and the target application, we also augment the model with new bases learned from the training set of the target application during the pruning process. This enables us to learn the new bases necessary for the target application that are absent in the pretrained model.
@@
We evaluate Basel on two models—Llama 2-7B and Llama 2-13B (Touvron et al., 2023b)—and two tasks—mathematical reasoning and code generation. We evaluate the pass@1 accuracy of the compressed models on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) for the mathematical reasoning task and HumanEval (Chen et al., 2021a) and MBPP (Austin et al., 2021) for the code generation task. Compared to state-of-the-art baselines, our approach achieves substantially better performance, improving accuracy by up to 19% when the compression ratio [^1] The compression ratio is defined as the ratio of the original model size to the compressed model size. exceeds 6 for mathematical reasoning and 4 for code generation. This also indicates that, in cases of deep compression, our method reduces model size by up to 2.7 times while maintaining comparable accuracy to baseline methods.
@@
  • We propose Basel, a low-rank decomposition approach to compress pretrained large language models for target applications. Basel identifies the beneficial and redundant components of large language models by relearning the importance (i.e., singular values) of bases using the training set of the target application, and then selects bases based on their importance.
  • We evaluate Basel across multiple tasks and models, demonstrating its superior performance in deep compression.
@@ 2 Related Work
Singular value decomposition (SVD) (Golub & Van Loan, 1996) has been applied to reduce the size of machine learning models. Prior research (Xue et al., 2013; Jaderberg et al., 2014; Denton et al., 2014; Zhang et al., 2015; Povey et al., 2018; Chen et al., 2018; Acharya et al., 2019; Noach & Goldberg, 2020) has developed various SVD algorithms to compress different components of models, such as DNNs, CNNs, and embedding layers for a range of applications including natural language processing, speech, and vision. The primary distinction between our work and these prior studies is that they do not relearn the importance of bases using the training data of target applications. Instead, they typically prune bases according to the singular values from the original or finetuned models. FWSVD (Hsu et al., 2022) evaluates the importance of individual weights rather than bases during SVD. As highlighted in Sections 1 and 3, the bases hold significant physical meanings. Considering importance at this level of granularity results in improved performance. As shown in Section 4, our approach surpasses FWSVD in deep compression performance.
A recent study (Sharma et al., 2024) applied SVD to large language models. Its focus is on determining the optimal rank for each layer, while our emphasis is on basis selection. The two methods are orthogonal but complementary and can be combined. Chen et al. (2021b) and Yu & Wu (2023) suggest reconstructing bases based on feature mimicking. These approaches are orthogonal to ours—they concentrate on basis reconstruction, whereas we focus on basis selection. Their methods complement ours and can be integrated together to achieve enhanced compression results.
3 Basel
In this section, we describe our proposed compression method, Basel.
For a linear layer $\mathbf{y}=\mathbf{W}\mathbf{x}+\mathbf{b}$, Singular Value Decomposition (SVD) factorizes its weight matrix $\mathbf{W}\in\mathbb{R}^{n\times m}$ as the product of three matrices $\mathbf{U}$, $\mathbf{S}$, and $\mathbf{V}$:
$\mathbf{W}=\mathbf{USV}^{\top}$ (1)
where $\mathbf{U}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\mathbf{S}=\mathrm{diag}\left(s_{1},\cdots,s_{r}\right)$, and $\mathbf{V}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$. The values $\left\{s_{i}\in\mathbb{R},i=1,\cdots,r\right\}$ are positive singular values. [^2] We drop zero singular values and the corresponding columns of matrices $U$ and $V$. The vectors $\left\{\mathbf{u_{i}}\in\mathbb{R}^{n},i=1,\cdots,r\right\}$ and $\left\{\mathbf{v_{i}}\in\mathbb{R}^{m},i=1,\cdots,r\right\}$ are orthonormal, i.e., $\left\|\mathbf{u_{i}}\right\|=1$, $\left\|\mathbf{v_{i}}\right\|=1$, $\mathbf{u_{i}}\perp\mathbf{u_{j}}$, and $\mathbf{v_{i}}\perp\mathbf{v_{j}}$ if $i\neq j$.
@@ 3 Basel
$\mathbf{W}=\sum_{i=1}^{r}s_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ (2)
Let matrix $\mathbf{W_{i}}=\mathbf{u}_{i}\mathbf{v}_{i}^{\mathrm{T}}$, then
$\left\|\mathbf{W_{i}}\right\|=\sqrt{\mathrm{tr}\left(\mathbf{W_{i}}^{\top}\mathbf{W_{i}}\right)}=\sqrt{\mathrm{tr}\left(\mathbf{v_{i}}\mathbf{u_{i}}^{\top}\mathbf{u_{i}}\mathbf{v_{i}}^{\top}\right)}=\sqrt{\mathrm{tr}(\mathbf{v}_{i}\mathbf{v_{i}}^{\mathrm{T}})}=\sqrt{\mathrm{tr}(\mathbf{v_{i}}^{\mathrm{T}}\mathbf{v_{i}})}=1$
$\left\langle\mathbf{W_{i},W_{j}}\right\rangle=\mathrm{tr}\left(\mathbf{W_{i}}^{\top}\mathbf{W_{j}}\right)=\mathrm{tr}\left(\mathbf{v_{i}}\mathbf{u_{i}}^{\top}\mathbf{u_{j}}\mathbf{v_{j}}^{\top}\right)=0,\quad\text{if}\;i\neq j$
Therefore, $\left\{\mathbf{u}_{i}\mathbf{v}_{i}^{\top},i=1,\cdots,r\right\}$ can be seen as a group of orthonormal bases in a subspace of $\mathbb{R}^{n\times m}$, and $\left\{s_{i},i=1,\cdots,r\right\}$ are their weights, making the weight matrix $\mathbf{W}$ a linear combination of these bases.
@@
$\mathbf{y}=\mathbf{W}\mathbf{x}+\mathbf{b}=\sum_{i=1}^{r}s_{i}\mathbf{u_{i}}\mathbf{v_{i}}^{\top}\mathbf{x}+\mathbf{b}=\sum_{i=1}^{r}s_{i}\left\langle\mathbf{x},\mathbf{v_{i}}\right\rangle\mathbf{u_{i}}+\mathbf{b}$
@@
In other words, for each basis (i.e., filter) $\mathbf{u}_{i}\mathbf{v_{i}}^{\top}$, the similarity between the input signal $\mathbf{x}$ and the unit direction vector $\mathbf{v_{i}}$ is measured by their inner product. This inner product is then multiplied by the (positive) singular value $s_{i}$ to determine the weight for the unit direction vector $\mathbf{u_{i}}$. The output signal $\mathbf{y}$ is the weighted sum of $\mathbf{u_{i}}$. Figure 2 illustrates this interpretation of the role of bases from the perspective of signal processing.
Рис. 2: An interpretation of the role of bases from the perspective of signal processing.
Large language models pretrained on diverse datasets learn bases that capture a wide range of semantic concepts. To examine this, we factorize the $\mathbf{W_{O}^{h}W_{V}^{h}}$ matrix from the attention layers of both the original Llama 2-7B model and its math-finetuned counterpart. Focusing on bases associated with large singular values, we extract the corresponding $\mathbf{u}$ vectors and interpret their semantics via a de-embedding process. Specifically, we project each $\mathbf{u}$ vector into the vocabulary space by passing it through the output layer (i.e., the lm_head), yielding a logit vector over the vocabulary. We then identify the token with the highest logit as the de-embedding result. Table 1 summarizes our findings. First, some bases encode highly specific semantic content, such as concepts related to Apple, Python, programming, locations, non-English characters, and math symbols. Second, even within the same layer, different bases can represent very different meanings. For example, in the math-finetuned model at layer 22, attention head 6, basis 1 corresponds to non-English characters, while basis 3 corresponds to math symbols. Third, fine-tuning does not necessarily remove irrelevant bases: the non-English character basis remains present even after fine-tuning on math-focused data.
Таблица 1: Meaning of bases in the vanilla and math-finetuned Llama 2-7B.
Domain Basis Top ten most probable tokens corresponding to the basis
Apple
Vanilla model, Layer 16,
Head 25, Basis 2
_iOS_Xcode_ios_Apple_Mac _iPhone_app_xcode_NS_App
Python
Vanilla model, Layer 17,
Head 6, Basis 1
._in<0x0A>_…_andĽ_tofor!
Location
Vanilla model, Layer 17,
Head 25, Basis 1
_Massachusetts_Illinois_Chicago_Boston_Dan _Harvard_Connecticut_IL_Bulg_Bulgar
Non-English
Math-finetuned model,
Layer 22, Head 6, Basis 1
:Yā’
Math
Math-finetuned model,
Layer 22, Head 6, Basis 3
[…]
These findings suggest that many bases in the model are useful for specific tasks, but may be irrelevant for others. When these irrelevant bases are used as filters in non-target applications, two scenarios can occur: the filter may not be activated (due to a small inner product $\left\langle\mathbf{x},\mathbf{v}_{i}\right\rangle$), or worse, the filter is activated, introducing harmful information into the output and degrading performance. This indicates that pruning such bases could reduce model size with minimal performance loss, and in some cases, even enhance performance for the target application.
@@
In our approach, Basel, we determine the importance of the bases from the pretrained model by retraining their singular values on the training set of the target application. The weight matrix $\widetilde{W}$ in Basel is represented as:
$\mathbf{\widetilde{W}}=\sum_{i=1}^{r}\tilde{{s}}_{i}\mathbf{u_{i}}\mathbf{v_{i}}^{\top}+\sum_{j=1}^{\tilde{r}}\tilde{\mathbf{u}}_{\mathbf{{j}}}\tilde{\mathbf{v}}_{\mathbf{{j}}}^{\top}$ (3)
@@
[…]
4 Convert the weight matrix of each linear layer in MM—excluding the embedding layer—into the form specified by equation equation 3;
@@
6  Tune the learnable parameters in equation equation 3;
7